# FreqKV: Frequency Domain Key-Value Compression for Efficient Context Window Extension

### v2

@@ 7 Conclusion
In this paper, we introduce FreqKV to compress KV states iteratively in the frequency domain for LLMs. We exploit the energy concentration of KV states in the frequency domain within the decoder layers. Specifically, we filter out the high-frequency components that are of low magnitude and retain the low-frequency components for compression. The KV cache is compressed in the frequency domain without introducing additional compression modules. Iteratively compressing the KV cache, FreqKV could extend the context window efficiently for LLMs. With minimal training of low-rank adaption, LLMs learn to leverage the compressed KV cache. Through extensive experiments and analysis on long context modeling and understanding, FreqKV demonstrates its efficiency and effectiveness in context extension.
Appendix A Head-wise Analysis of Power Spectrums
We explore the power spectrum distribution of key states and value states in different attention heads of LLaMA-2-7b as shown in Figure 6. Although values of power spectrums vary in different heads, their distribution exhibits similar patterns. They have a consistent tendency to aggregate energy in the low-frequency components along the decoding procedure. It could be promising to study specific differences and associations in different heads or other modules.
@@ Appendix A Head-wise Analysis of Power Spectrums
Appendix B Compressing Key States with RoPE
@@
We train LLaMA-2-7b on RedPajama with the variant method that key states are compressed after applying RoPE. The training length is 8K. Models are evaluated on the test set of PG-19 and Proof-pile. The results in Table 4 demonstrate that the variant suffers from performance degradation, which is probably due to the usage of position embeddings out of the original window. In contrast, FreqKV compresses and caches key states before applying RoPE. Therefore, they are allocated with position indices within the original window rather than the lengthy sequence, which results in the context window extension with no need of using out-of-bound position embeddings.
@@
Appendix C Training settings
@@
For long context language modeling, we fine-tune LLaMA-2-base on the RedPajama (Computer,, 2023) pre-training dataset for 1000 steps. For long context understanding, we SFT LLaMA-2-chat and LLaMA-3-instruct on the instruction following dataset LongAlpaca (Chen et al.,, 2024) for 5 epochs, which consists of 6.28K long-context QA samples.
@@
The total batch size ($GPU\_number\times Batch\_size\_per\_device\times Gradient\_accumulation\_steps$) is 64. The learning rate increases linearly from 1e-6 to 2e-5 with 20 warm-up steps and remains constant in the following steps. The rank used in the LoRA (Hu et al.,, 2021) fine-tuning is set to 8. Following LongLoRA (Chen et al.,, 2024), the embedding and normalization layers are learnable during training. All the experiments are conducted on ADA6000 and H100 GPUs. Moreover, we equip our method with FlashAttention-2 (Dao,, 2023) for further acceleration and memory saving.
The training time and memory requirements of FreqKV and Full FT on the RedPajama are reported in Table 5. For comparison, we set $Batch\_size\_per\_device$ to 1 and $Gradient\_accumulation\_steps$ to 8. 8$\times$×ADA6000-48GB GPUs are used. Statistics show that FreqKV significantly reduces both training time and memory requirements, demonstrating its efficiency.
Methods Training Length Time (hours) Memory (GB)
Full FT 8K 28.88 44.04
FreqKV 8K 17.07 24.58
16K 38.57 34.00
32K 89.35 45.37
Appendix D Compression Overhead
@@
We conduct further studies on the compression overhead of FreqKV. As introduced in Section 4.3, chunk-wise attention is performed for the prefilling tokens. The size of the attention matrix in each chunk is $(N-L-S)\cdot N$ except for the last few tokens. Therefore, the computational cost of self-attention grows approximately linearly with the input length like sliding window attention (Beltagy et al.,, 2020).
@@
We use torchprofile [^1] https://github.com/zhijian-liu/torchprofile to count the number of Floating Point Operations (FLOPs) with input sequences of different lengths for LLaMA-2-7b. We calculate FLOPs for the model with the original attention which leverages full KV states, and with FreqKV with the retaining ratio of 0.5. To quantify the compression overhead, we have measured FLOPs of the Dropping, which retains the most recent tokens instead of compressing in the frequency domain. It shares the same sink size and retaining size as FreqKV. The difference in FLOPs between the two methods shows the overhead of compression.
@@
The statistics are given in Table 6. The compression times of FreqKV with different context lengths are also reported in the table. It shows that the computation overhead of our compression process grows less than 0.5% even with a length of 32K, which could be negligible. FreqKV reduces more FLOPs as the input length grows from 4K to 32K compared to Full KV. This is because the compression is performed every $N-L-S$ tokens with the complexity of $O(N\text{log}N)$, which is negligible compared to the quadratic self-attention.
@@
Table 6: FLOPs (TFLOPs) with input sequences of different lengths. Experiments are conducted on the ADA6000 GPU of 48GB, where full KV of 12K tokens with float16 will cause an OOM (Out-of-Memory) issue. The difference between FreqKV and Dropping shows the computation overhead of compression.

### v3

@@ 6 Conclusion
In this paper, we introduce FreqKV, a frequency-domain KV cache compression framework that exploits the energy concentration of KV states to retain low-frequency components while discarding redundant high-frequency ones for LLMs. By iteratively compressing the KV cache, FreqKV achieves efficient long-context extension while preserving decoding quality. Without introducing additional parameters or architectural modifications, FreqKV exhibits remarkable efficiency and effectiveness on both prefilling and decoding stages. Extensive experiments across language modeling, generation, and understanding benchmarks demonstrate that FreqKV not only consistently surpasses existing KV compression methods but also offers a viable and efficient approach for context window extension.
Ethics Statements
Our work pertains to key-value compression and context extension of large language models. In this work, we use only publicly available data and artifacts. There are no ethical issues in our paper, including its motivation and experiments.
Reproducibility Statements
We have provided detailed implementations of our method throughout the paper. Our method is comprehensively elaborated in Section 3. Detailed settings of our experiments and analyses are given in Section 4,  5, and Appendix F. Our code is included in the Supplementary Material for reproducibility.
The Use of Large Language Models
We used ChatGPT solely as a language editing tool to polish grammar, improve clarity, and refine the academic style of the manuscript. All research ideas, methods, experiments, and analyses were independently developed and conducted by the authors.
Appendix A Discrete Cosine Transform
[…]
where $t=0,1,\cdots,N-1$. $\alpha_{t}$ is the normalization factor. The time-domain signal ${\bm{X}}_{0:N-1}$ can be recovered by applying the inverse DCT (IDCT) on the frequency components ${\bm{Z}}_{0:N-1}$:
$x_{n}=\sum_{t=0}^{N-1}\alpha_{t}\cdot z_{t}\cdot\cos\left[\frac{\pi t(2n+1)}{2N}\right].$ (7)
The frequency components are expressed as a combination of the original signals. The values can be computed using the Fast Fourier Transform (FFT) with a complexity of $O(N\text{log}N)$. The amplitudes of frequency components are utilized in the power spectrum analysis to represent the energy or magnitude of components. The components of higher energy in the frequency domain indicate they are more informative (He et al., 2023).
For time signals (e.g., texts) and space signals (e.g., images), their energy is most concentrated in low-frequency DCT components. This phenomenon occurs due to the mathematical properties of DCT and the inherent smoothness of real-world signals. The DCT expresses a signal as a sum of cosine functions oscillating at different frequencies. In the case of text embeddings, neighboring tokens exhibit strong temporal correlations, meaning adjacent values in the semantic space tend to be similar. Low-frequency cosine components are slowly varying and capture these smooth variations well, while high-frequency components mainly represent abrupt changes or noise.
@@ Appendix A Discrete Cosine Transform
Appendix B Head-wise Analysis of Power Spectrums
We explore the power spectrum distribution of key states and value states in different attention heads of LLaMA-2-7B, as shown in Figure 5. Although the values of power spectrums vary in different heads, their distribution exhibits similar patterns. They have a consistent tendency to aggregate energy in the low-frequency components along the decoding procedure. It could be promising to study specific differences and associations in different heads or other modules.
Appendix C Compressing Key States with RoPE
@@
We train LLaMA-2-7B on RedPajama with the variant method that key states are compressed after applying RoPE. The training length is 8K. Models are evaluated on the test set of PG-19 and Proof-pile. The results in Table 8 demonstrate that the variant suffers from performance degradation, which is probably due to the usage of position embeddings out of the original window. In contrast, FreqKV compresses and caches key states before applying RoPE. Therefore, they are allocated with position indices within the original window rather than the lengthy sequence, which enables context window extension with no need of using out-of-bound position embeddings.
@@
Appendix D Training settings
@@
For long context language modeling, we fine-tune LLaMA-2-base on the RedPajama (Computer, 2023) pre-training dataset for 1000 steps. For long context understanding, we SFT LLaMA-2-chat and LLaMA-3-instruct on the instruction following dataset LongAlpaca (Chen et al., 2024) for 5 epochs, which consists of 6.28K long-context QA samples.
@@
The total batch size ($GPU\_number\times Batch\_size\_per\_device\times Gradient\_accumulation\_steps$) is 64. The learning rate increases linearly from 1e-6 to 2e-5 with 20 warm-up steps and remains constant in the following steps. The rank used in the LoRA (Hu et al., 2021) fine-tuning is set to 8. Following LongLoRA (Chen et al., 2024), the embedding and normalization layers are learnable during training. Moreover, we equip our method with FlashAttention-2 (Dao, 2023) for further acceleration and memory saving.
Table 9: Perplexity evaluation on the test set of Proof-pile. The superscript “*” means that we reproduce LoCoCo following their official code for evaluation. The results of full fine-tuning and LongLoRA are reported from LongLoRA (Chen et al., 2024). Full FT and LoCoCo encounter the OOM (Out-of-Memory) issue with the training length of 16K.
Size Training Length Method Inference Cache Evaluation Context Length
2048 4096 8192 16384 32768
7B 8192 Full FT Full 3.14 2.85 2.66 - -
LoCoCo Compressed 3.40 3.20 2.88 - -
LongLoRA Full 3.20 2.91 2.72 - -
FreqKV Compressed 3.15 2.88 2.79 2.76 2.75
16384 Full FT Full OOM during Training
LoCoCo Compressed OOM during Training
LongLoRA Full 3.17 2.87 2.66 2.51 -
FreqKV Compressed 3.15 2.88 2.78 2.74 2.73
32768 LongLoRA Full 3.35 3.01 2.78 2.61 2.50
FreqKV Compressed 3.16 2.89 2.78 2.74 2.72
13B 8192 Full FT Full 2.96 2.69 2.53 - -
LongLoRA Full 3.04 2.77 2.60 - -
FreqKV Compressed 2.99 2.74 2.65 2.63 2.63
Appendix E Evaluation on Proof-pile
In Table 9, we present the evaluation results on the Arxiv math dataset Proof-pile. On Proof-pile, FreqKV trails LongLoRA at longer context lengths. This gap arises because Proof-pile, unlike natural language corpora, contains dense symbolic and mathematical expressions where minor variations (e.g., a single symbol or index) drastically alter semantics. Such localized variations are more difficult to preserve under compression, making it particularly challenging for models using compressed cache to predict the next token.
LongLoRA retains the full KV cache during inference, avoiding any information loss and thus performing better on formula-heavy tasks. Nonetheless, FreqKV remains competitive and notably outperforms another compression-based method, LoCoCo. Additionally, the performance of FreqKV does not degrade further as the context length increases, highlighting its robustness.
Appendix F Ablation Study
F.1 Retaining High/Low-Frequency Components
We further compare variants of FreqKV that retain only high-frequency vs. only low-frequency components using LLaMA-2-base-7B on the validation set of PG-19. Results (PPL) are provided in Table 10.
It shows that low-frequency components adapt better than the variant that retains high-frequency parts, which is consistent with our analysis in summarization tasks (Table 1). While the variant deteriorates when exceeding the original context window, low-frequency components, which carry more information, benefit the model, and the perplexity remains stable. This result aligns well with Frequency Principle (Xu et al., 2019; Luo et al., 2019) that low frequencies converge faster than high frequencies for Deep Neural Networks.
Table 10: Perplexity evaluation on PG-19 with high/low-frequency components retained.
Components Retained Evaluation Context Length
2048 4096 8192 16384
high-frequency 8.29 7.92 8.42 8.70
low-frequency 8.20 7.84 7.74 7.73
Figure 7: Perplexity (Line) and decoding latency (Bar) with different retaining ratios.
F.2 Retaining Ratio
We train LLaMA-2-7B on RedPajama (Computer, 2023) with different retaining ratios. The training length is 8K. Models are evaluated on the validation set of PG-19 with the evaluation length ranging from 2K to 64K.
The performance of FreqKV with different retaining ratios is presented in Figure 8. The decoding time (seconds) is averaged over five runs. FreqKV performs better with more frequency components retained. However, larger retaining ratios lead to more frequent compression steps and denser attention computation, hence higher latency.
We also evaluate the performance of FreqKV when modifying its retaining ratio during evaluation on the validation set of PG-19. The PPL scores at a length of 8K are presented in Table 11. These results suggest that FreqKV is robust to mismatched training/evaluation retaining ratios, especially when using larger inference ratios than training (which retains more information).
Table 11: Perplexity evaluation on PG-19 with mismatched training/evaluation retaining ratios.
Retaining Ratio during Training Retaining Ratio during Evaluation
0.25 0.5 0.75
0.25 7.84 7.81 7.82
0.5 7.89 7.76 7.77
0.75 7.97 7.79 7.75
F.3 Sink Size
The effect of sink token count has been explored in StreamingLLM (Xiao et al., 2024b). We also ablate the effect of the sink token number $S$ for FreqKV on LLaMA-2-7B. Results are given in Table 12. It shows that a threshold of four sink tokens appears to be enough, with subsequent additions contributing marginal effects.
Table 12: Perplexity evaluation on PG-19 with different numbers of sink tokens.
Sink Size Evaluation Context Length
2048 4096 8192
0 8.23 7.87 7.84
2 8.21 7.85 7.78
4 8.20 7.84 7.76
8 8.20 7.83 7.75
F.4 Effect of Frequency-Domain Compression
(a) Dropping.
(b) FreqKV.
Figure 9: Comparison of FreqKV and its variant Dropping on LLaMA-2-chat-7B.
To demonstrate the effectiveness of our compression operation, a variant method, Dropping, is also implemented and evaluated on Needle-in-a-Haystack for comparison, as in Figure 9. Instead of compressing in the frequency domain, Dropping retains the most recent tokens like StreamingLLM (Xiao et al., 2024b). Although trained under identical conditions, Dropping suffers from an irreversible loss of early-context information. FreqKV, by contrast, preserves essential signals and achieves superior performance.
F.5 Effect of rescaling
Figure 10: Curves of training loss for FreqKV with and without rescaling.
In Equation 3, $\sqrt{\frac{L}{N}}=\sqrt{\gamma}$ works as a rescaling factor for the compressed signals to restore the original amplitude when conducting IDCT. In DCT, an $N$-length time-domain signal is normalized by $\sqrt{N}$; during IDCT, the retained $L$ frequency components are correspondingly normalized by $\sqrt{L}$. When iterative compression is applied across multiple steps, this mismatch in normalization scales can cause the reconstructed signals to be progressively amplified. In our pilot study, the training loss is significantly higher at the early stages when the compressed signals are not rescaled as presented in Figure 10. Therefore, we introduce the rescaling factor $\sqrt{\frac{L}{N}}$ to restore the original amplitude in the time domain.
Appendix G Compression Overhead
We conduct further studies on the compression overhead of FreqKV. FreqKV performs compressions on-the-fly during decoding. We compare the decoding time required by LLaMA-2-7B with the full cache and our FreqKV when the sequence length increases. As shown in Figure 8, the decoding time starts to diverge at the length of 4K. While the full cache utilization leads to a quadratic growth in decoding time, the decoding time of FreqKV increases approximately linearly, with a negligible time spent on compression, showcasing its efficiency.
@@ Appendix G Compression Overhead
As introduced in Section 3.4, chunk-wise attention is performed for the prefilling tokens. The size of the attention matrix in each chunk is $(N-L-S)\cdot N$ except for the last few tokens. Therefore, the computational cost of self-attention grows approximately linearly with the input length, like sliding window attention (Beltagy et al., 2020).
@@
We use torchprofile [^1] https://github.com/zhijian-liu/torchprofile to count the number of Floating Point Operations (FLOPs) with input sequences of different lengths for LLaMA-2-7B. We calculate FLOPs for the model with the original attention, which leverages full KV states, and with FreqKV with the retaining ratio of 0.5. To quantify the compression overhead, we have measured FLOPs of Dropping, which retains the most recent tokens instead of compressing in the frequency domain. It shares the same sink size and retaining size as FreqKV. The difference in FLOPs between the two methods shows the overhead of compression.
@@
The statistics are given in Table 13. The compression times of FreqKV with different context lengths are also reported in the table. It shows that the computation overhead of our compression process grows less than 0.5% even with a length of 32K, which could be negligible. FreqKV reduces more FLOPs as the input length grows from 4K to 32K compared to Full KV. This is because the compression is performed every NLSN-L-S tokens with the complexity of $O(N\text{log}N)$, which is negligible compared to the quadratic self-attention.
@@
Table 13: FLOPs (TFLOPs) with input sequences of different lengths. Experiments are conducted on a single NVIDIA RTX6000 Ada-48GB GPU, where full KV cache of 12K tokens with float16 will cause an OOM (Out-of-Memory) issue. The difference between FreqKV and Dropping shows the computation overhead of compression.